OpenStudy (anonymous):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π(x - 7)2 (x - 3) What does the factor π (x - 7)2 represent? The total surface area of the tank The area of the curved sides of the tank The area of the circular base of the tank The area of the base of about three such tanks

OpenStudy (jdoe0001):

"height of the tank is (x - 3)" -> h = (x-3) "radius of a cone-shaped tank is 4 feet less than its height." -> r = (x-3) - 4 \(\bf \textit{volume of a cone}=\cfrac{\pi {\color{red}{ r}}^2 {\color{blue}{ h}}}{3}\qquad r={\color{red}{ (x-3)-4}}\qquad h ={\color{blue}{ (x-3)}} \)

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

So it would be C ?

OpenStudy (jdoe0001):

well... not quite.... do you know what the area of a circle is?

OpenStudy (anonymous):

3.14 right ?

OpenStudy (anonymous):

pie...

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

\(\bf \textit{volume of a cone}=\cfrac{\pi {\color{red}{ r}}^2 {\color{blue}{ h}}}{3}\qquad r={\color{red}{ (x-3)-4}}\qquad h ={\color{blue}{ (x-3)}} \\ \quad \\ \implies \textit{volume of this cone}=\cfrac{\pi {\color{red}{ [(x-3)-4]}}^2 {\color{blue}{ (x-3)}}}{3} \\ \quad \\ \textit{keep in mind that }\textit{area of a circle }=\pi r^2\)

OpenStudy (anonymous):

Ohhhh okay so it would be a the total surface area of the tank ?

OpenStudy (jdoe0001):

is it?

OpenStudy (anonymous):

Thats what makes the most sense to me...

OpenStudy (jdoe0001):

well.. if you were to simplify (x-3) -4 what would you get?

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

wouldnt it be like 1x

OpenStudy (jdoe0001):

you may want to review your linear simplification

OpenStudy (jdoe0001):

(x-3) - 4 = x -3 - 4